Two forces act on a 17.2 kg object, A= 27N @27° and B= 43N @ x°. The direction of B is __ degrees if the acceleration is 3.79477 m/s^2 @ 0°
1
Expert's answer
2020-12-14T07:16:03-0500
"F=ma=(17.2)(3.79477)=65.27\ N"

"F=27\cos{27}+43\cos{x}\\
x=\arccos{\frac{65.27-27\cos{27}}{43}}=16.58\degree"